cathrinerose23:

5,000 mililiters is what in liters?

Luigi0210:

there are 1000 mL in 1 L Sooo \(\Large \frac{ 5000 mL}{1 } * \frac{1 L}{1000 mL} = \frac{5000}{1000} = ? Liters\)

cathrinerose23:

ok

cathrinerose23:

its 5 thx bro

Luigi0210:

No prob

cathrinerose23:

lol